OpenStudy (anonymous):

The frequency of G5 is 783.99 Hz. What is the frequency of the note that is a perfect fifth above G5?

OpenStudy (wolf1728):

I believe that musical notes are based on the 12th root of 2 which is 1.0594631 So if G5 is 783.99 Hz then G#5 would be 783.99 Hz * 1.0594631 = 830.61 A5 would be 783.99 * (1.0594631)^2 = 880.00 I think that should help you out a LOT.

OpenStudy (wolf1728):

I believe a perfect 5th above G would be D correct? The notes to get there are G#, A, Bb, B, C, C#, D So you would get 783.99 and multiply it by 1.0594631 to the 7th power.

OpenStudy (wolf1728):

Need more help? Okay 783.99 * (1.0594631)^7 = 783.99 * 1.49830707688 = 1,174.66 Hz Okay now?

OpenStudy (wolf1728):

By the way I am a musician. I play electric guitar and love rock and roll. However, just because it is rock and roll, some folks actually want to know about the physics of music.

OpenStudy (wolf1728):

Gee. katrina you aren't very talkative. Anyway, I'm glad to have helped you out. :-)